28.0g Na_2S_2O_8 Express your answer with the app
Answer #1

No. of moles = Weight of the substance / molecular mass

1) Weight of Na2S2O8 = 28.0 grams

Molecular mass of Na2S2O8 = 238.09 gram/mol

no. of moles = 28.0 / 238.09 = 0.1176 moles = 11.76 x 10-2 moles

2)

Weight of ZnF2 = 14.9 grams

Molecular mass of ZnF2 = 103.377 gram/mol

no. of moles = 14.9 / 103.377 = 0.14413 moles = 14.413 x 10-2 moles